An enzyme kinetics experiment is performed with the enzyme Imaginase, which converts Horses into Unicorns with a known Kcat of 300 s-. In an experiment with (Imaginase) at 40 nM and (Horses) at 10 uM, the researchers find that Vo = 4 uM/s. What is the Km of Imaginase for its substrate Horses?
Answer #1

Keat=300/= Vmax Et (S)=10um (Et) = 400 M = 40 -0.04 UM Vo= 4 UMIS km=? - Vo=Vmax (8) tnt(s) 300 - Vmax - Vmax= 300X0604 0004

Please give me a good rating.